OpenStudy (anonymous):

Can this be solved w/o graphing? sin 3x = cos x

OpenStudy (anonymous):

I think you will need to use triple angle identities

OpenStudy (anonymous):

you use the double angle formula or half angle formula

myininaya (myininaya):

i dont know if it possible to find exact solutions

OpenStudy (anonymous):

it is because because you haave to play around with it to find the correct answer it sums up to equal the other side

myininaya (myininaya):

oh i got it

myininaya (myininaya):

we have sin3x=sin(pi/2-x) 3x=pi/2-x 4x=pi/2 x=pi/8

OpenStudy (anonymous):

just keep trying

myininaya (myininaya):

i dont have to i got it lol

OpenStudy (anonymous):

well congratz

myininaya (myininaya):

we call also make a answer general by saying x=pi/8+2npi for n=..-2,-1,0,1,2,...

OpenStudy (anonymous):

yu can say that but stil make sure that is definately the right answer its alwayz good to double check

myininaya (myininaya):

and we also have x=5pi/8+2npi for n=..-2,-1,0,1,2,..

OpenStudy (nikvist):

\[x=\frac{\pi}{4}+\pi k\quad\vee\quad x=\frac{\pi}{8}+\frac{\pi}{2}k\quad,\quad k\in\mathbb{Z}\]